Circular No. 8045

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003E, 2003F, 2003G, 2003H Discoveries of three supernovae have been found on unfiltered CCD

images and reported to the Central Bureau. All three objects were found by J. Graham and W. Li (via LOTOSS/KAIT; cf. IAUC 7906); SN 2003F was found independently by T. Puckett, Mountain Town, GA, with D. Toth (Lima, OH) via the Puckett Observatory 0.50-m MassImaging reflector (confirmed Jan. 9.13; Puckett’s data for SN 2003F tabulated below).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003F Jan. 8.16 5^{h}11^{m}32.^{s}99 +17^{o}03′28′′.4 16.7 7′′.3 E, 6′′.1 N
2003G Jan. 8.2  2 08 28.13 + 6 23 51.9 16.2 6′′.0 E, 9′′.8 N
2003H Jan. 8.3  6 16 25.68 −21 22 23.8 17.8 31′′.9 W, 9′′.6 N

The offsets of SN 2003H are relative to the eastern component of the interacting galaxy pair. Li provides position end figures 33^{s}.03, 28′′.4 for SN 2003F. Additional magnitudes [sources: ‘K’ = KAIT; ‘S’ = Palomar Sky Survey; ‘P’ = Puckett; ‘Q’ = Y. L. Qiu, Beijing Astronomical Observatory (via Li)]: SN 2003F in UGC 3261, 1953 Jan. 9, [20.0: (S, P); 1983 Sept. 9, [21.0: (S, P); 2002 Feb. 14, [20.5: (P); Dec. 11, [19.0 (K); 2003 Jan. 8.3, 16.3 (K); 9.6, 16.1 (Q). SN 2003G in IC 208, 2002 Dec. 12, [19.0 (K); 2003 Jan. 9.5, 15.8 (Q). SN 2003H in NGC 2207, Jan. 3, [19.0 (K); 9.6, 18.0 (Q).

M. Hamuy, Carnegie Observatories, reports that spectroscopic observations (range 320–920 nm), made by J. Maza (University of Chile) on Jan. 9.11 UT with the Las Campanas Observatory Baade 6.5-m telescope (+ Boller & Chivens spectrograph), show that SN 2003F is a type-Ia supernova, very similar to SN 1994D at 3 days before maximum. The minimum of the Si II 635.5-nm line yields an expansion velocity of 10400 km/s. The interstellar Na I D line is present in absorption with an equivalent width of 0.16 nm, and this suggests some reddening in the host galaxy. A spectrum of SN 2003G, from Jan. 9.05, has a blue continuum and a strong Hα line in emission with an FWHM of 400 km/s, superposed on a broader emission component, indicating that the supernova is of type IIn; a narrow absorption is present in the blue wing of Hα, and the spectrum bears strong resemblance to that of SN 1994W (Filippenko 1997, ARAA 35, 309).

Hamuy adds that a spectrum of SN 2003E (cf. IAUC 8044), obtained as above on Jan. 9.15, shows it to be a type-II supernova with a blue continuum and Balmer lines with prominent P-Cyg profiles. The minimum of Hβ yields an expansion velocity of 8200 km/s.

2003 January 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green